OpenStudy (anonymous):

A tour group split into two groups. First group bought 6 slices of pizza and 4 drinks for $28.28. Second group bought 5 slices and 4 drinks for $24.50. How much is one slice of pizza?

OpenStudy (anonymous):

I get to 6x+4y=28.28 5x+4y+24.50 Then I get stuck...

OpenStudy (anonymous):

k u got the equations down now do (6x+4y=28.28)-(5x+4y=24.50)

OpenStudy (anonymous):

u should get x=3.78

OpenStudy (anonymous):

thank you. Just having you write that out helped me solve it.

OpenStudy (anonymous):

no prob^^